# Towards Assessing Data Replication in Music Generation with Music Similarity Metrics on Raw Audio

### v1

@@ Abstract
Recent advancements in music generation are raising multiple concerns about the implications of AI in creative music processes, current business models and impacts related to intellectual property management. A relevant challenge is the potential replication and plagiarism of the training set in AI-generated music, which could lead to misuse of data and intellectual property rights violations. To tackle this issue, we present the Music Replication Assessment (MiRA) tool: a model-independent open evaluation method based on diverse audio music similarity metrics to assess data replication of the training set. We evaluate the ability of five metrics to identify exact replication, by conducting a controlled replication experiment in different music genres based on synthetic samples. Our results show that the proposed methodology can estimate exact data replication with a proportion higher than $10\%$. By introducing the MiRA tool, we intend to encourage the open evaluation of music generative models by researchers, developers and users concerning data replication, highlighting the importance of ethical, social, legal and economic consequences of generative AI in the music domain. Code and examples are available for reproducibility purposes [^1] https://github.com/roserbatlleroca/mira.
@@ 1 Introduction
Significant advancements in generative algorithms for digital art creation are challenging the role of artificial intelligence (AI) in artistic practices. Regarding generative AI in music, there is an increasing discussion related to the use of computational tools in music creative processes [1], the effects on artists’ work, existing listening experiences and business models, and the impacts on intellectual property (IP) management [2, 3]. A key challenge is the potential replication and plagiarism of the training set in AI-generated music [3, 4], which can lead to data misuse and intellectual property violations.
The inherent opaque nature of music generation models challenges tracing references in the training set used in AI-generated music, limiting interpretation of whether generated samples contain replicated fragments. In addition, diffusion models, one of the most popular generative AI architectures, tend to memorise and replicate training data [5, 6, 7]. Understanding the behaviour of these models has become critical to address legal issues [8], especially when dealing with data protected by IP rights. This is significant in the music domain as the vast majority of music is protected by authorship and copyright.
@@
Thus, this work proposes assessing the effectiveness of five music similarity metrics [^2] Hereafter, music similarity metrics refer to audio-based metrics. (four standard widely-used and a novel one) in estimating exact data replication in music. We review the implications of potential data replication in AI-generated music (Section 2) and present our experimental setup, including the selected music similarity metrics and specific methodology to control and estimate exact data replication (Section 3). We analyse metrics’ behaviour in different music materials (Section 4.1), aiming to assess later their data replication detection sensitivity (Section 4.2). The proposed methodology is implemented in tool MiRA (Music Replication Assessment), which computes music similarity between reference and target samples to obtain global and per-pair distances (Section 5). Finally, we discuss our research’s insights, limitations and future perspectives (Section 6).
By introducing the MiRA tool, we advance towards the assessment of data replication in AI-generated music using similarity metrics, contributing to open evaluation methods for researchers, developers and users. We strive to raise awareness, detect and prevent misappropriation of training sets, and hope to motivate research on these issues.
@@ 2.1 Implications of potential data replication in AI-generated music
Music-generative AI is advancing rapidly with novel high-quality models driven by a strong push from the industry, which is encouraging a suitable environment for real-world deployment. Yet, music generation algorithms bring significant concerns regarding their ethical, social, legal and economic implications. A key challenge is the potential data replication in AI-generated music—inquiring whether a generative model extracts and copies fragments from the training data and whether AI-generated music can be considered novel and original [3, 4]. This issue is further complicated by the implications derived concerning data misuse and IP violations such as copyright infringement. Moreover, diffusion models, one of the most popular architectures for generative AI, present high risks of data replication as they tend to memorize their training data [5, 6, 7]. In the image generation domain, Somepalli et al. [9] demonstrate instances where generated images with diffusion models contain object-level copies of their training data. Based on image retrieval frameworks, they compare generated images with training samples and detect when content has been replicated. Similarly, Carlini et al. [5] demonstrate that diffusion models memorize and reproduce images from their training data.
@@ 3.1 Audio Music Similarity Metrics
For this study, we consider five music similarly metrics: four standard and a novel one, covering a diversity of characteristics, from audio embeddings to state-of-the-art metrics. We here describe the metrics (summarised in Table 1) and methods used to implement them [^3] Two of the metrics rely on Essentia implementation. Essentia is an open-source library and tools for audio and music analysis, description and synthesis, developed in the Music Technology Group at Universitat Pompeu Fabra: https://essentia.upf.edu..
Cover Song Identification (CoverID) [19, 20, 21]: Cover song identification is a task aiming to detect whether two music recordings are based on the same composition, accounting for variations in tempo, structure, and instrumentation while keeping a similar melodic or harmonic line. CoverID relies on pitch-content features and local alignment. To obtain CoverID distance, we use the implementation available in Essentia [^4] https://essentia.upf.edu/reference/std_CoverSongSimilarity.html. A low CoverID value suggests substantial composition similarity between the two analysed music samples.
@@
Table 1: Summary of the considered music similarity metrics in this study, indicating their similarity trend ($\downarrow$/$\uparrow$).
@@ 3.2 Experimental Approach
For this experiment, we use an in-house dataset of 30-second audio previews from the Spotify API [^10] https://developer.spotify.com/documentation/web-api, composed of over 18,000 samples and 24 music genre classes. We focus on six music genre classes defined by Spotify API internal class labels: heavy metal, afrobeats, techno, dub, cumbia and bolero. These genres were chosen for their diverse musical compositions and elements, allowing us to examine the metrics across multiple scenarios. This selection aligns with ChatGPT’s recommendation to include genres with distinct musical characteristics.
@@ 4.1 Analysing metric behaviour
Figures 2, 3, 4, 5 and 6 depict the average $\mu$ and standard deviations $\sigma$ of the different metrics per degree of replication and music genre. We observe a steady and similar behaviour by three metrics (CoverID, CLAP and DEfNet) through all studied music genres, showing higher similarity values for cases with higher replication levels ($50\%$). Standard deviation decreases with increasing replication level, which suggests less disparity within the analysed pairs. These three metrics show the sensitivity [^11] Sensitivity is understood as the capability to differentiate between degrees of replication. to estimate data replication. Instead, KL divergence presents a surprising behaviour with very similar values of $\mu$ and $\sigma$ for different replication levels. We also observe a certain degree of sensitivity in all music genres, except for dub, where the baseline mean $\mu_{b}$ is smaller than in replication cases $\mu_{r}$, despite the standard deviation being higher ($\mu_{b}$=0.757, $\sigma_{b}$=0.511; $\mu_{r}$=0.862, $\sigma_{r}$=0.462). KL divergence demonstrates the capability of detecting replication but is ineffective in distinguishing between degrees of replication.
@@ 6 Discussion and Conclusions
This investigation focused on validating the use of music similarity metrics for assessing data replication in AI-generated music. We hypothesise that similarity metrics are effective in estimating data replication. Therefore, we framed the scope of our study to exact data replication in music samples, while conducting a controlled forced-replication experiment with synthetic data.

### v2

@@ Abstract
Recent advancements in music generation are raising multiple concerns about the implications of AI in creative music processes, current business models and impacts related to intellectual property management. A relevant discussion and related technical challenge is the potential replication and plagiarism of the training set in AI-generated music, which could lead to misuse of data and intellectual property rights violations. To tackle this issue, we present the Music Replication Assessment (MiRA) tool: a model-independent open evaluation method based on diverse audio music similarity metrics to assess data replication. We evaluate the ability of five metrics to identify exact replication by conducting a controlled replication experiment in different music genres using synthetic samples. Our results show that the proposed methodology can estimate exact data replication with a proportion higher than $10\%$. By introducing the MiRA tool, we intend to encourage the open evaluation of music-generative models by researchers, developers, and users concerning data replication, highlighting the importance of the ethical, social, legal, and economic consequences. Code and examples are available for reproducibility purposes. [^1] https://github.com/roserbatlleroca/mira
@@ 1 Introduction
Significant advancements in generative algorithms for digital art creation are challenging the role of artificial intelligence (AI) in artistic practices. Regarding generative AI in the music domain, there is an increasing discussion related to the use of computational tools in music creative processes [1], the effects on artists’ work, existing listening experiences and business models, and the impacts on intellectual property (IP) management [2, 3]. A key point is the potential replication and plagiarism of the training set in AI-generated music [3, 4], which can lead to data misuse and IP violations.
The inherent opaque nature of music generation models complicates tracing replications or references to training set samples in AI-generated music, limiting the interpretation of whether generated samples contain replicated fragments. In addition, diffusion models, one of the most popular generative AI architectures, tend to memorise and replicate training data [5, 6, 7]. Understanding the behaviour of these models has become critical to address legal issues [8], especially when dealing with data protected by IP rights. This is significant in the music domain as the vast majority of music is protected by authorship and copyright.
@@
Thus, this work proposes assessing the effectiveness of five music similarity metrics [^2] Hereafter, music similarity metrics refer to audio-based metrics. (four state-of-the-art widely-used and a novel one) in estimating exact data replication in music. We review the implications of potential data replication in AI-generated music (Section 2) and present our experimental setup, including the selected music similarity metrics and specific methodology to control and estimate exact data replication (Section 3). We analyse metrics’ behaviour in different music materials (Section 4.1), aiming to assess later their data replication detection sensitivity (Section 4.2). The proposed methodology is implemented in tool MiRA (Music Replication Assessment), which computes music similarity between reference and target samples to obtain global and per-pair distances (Section 5). Finally, we discuss our research’s insights, limitations and future perspectives (Section 6).
By introducing MiRA tool, we advance towards the assessment of data replication in AI-generated music using similarity metrics, contributing to open evaluation methods for accessibility for researchers, developers and users. We strive to raise awareness, detect and prevent misappropriation of training sets, and hope to motivate research on these issues.
@@ 2.1 Implications of potential data replication in AI-generated music
Music-generative AI is advancing rapidly with novel high-quality models driven by a strong push from the industry, which is encouraging a suitable environment for real-world deployment. Yet, music generation algorithms bring significant concerns regarding their ethical, social, legal and economic implications. A key challenge is the potential data replication in AI-generated music—inquiring whether a generative model extracts and copies fragments from the training data and whether AI-generated music can be considered novel and original [3, 4]. This issue is further complicated by the implications derived concerning data misuse and IP violations such as copyright infringement. Moreover, diffusion models, one of the most popular architectures for generative AI, present high risks of data replication as they have shown a tendency to memorise their training data [5, 6, 7]. In the image generation domain, Somepalli et al. [9] demonstrate instances where generated images with diffusion models contain object-level copies of their training data. Based on image retrieval frameworks, they compare generated images with training samples and detect when content has been replicated. Similarly, Carlini et al. [5] demonstrate that diffusion models memorize and reproduce images from their training data.
@@ 3.1 Audio Music Similarity Metrics
For this study, we consider five music similarly metrics: four state-of-the-art approaches and a novel one, covering a diversity of characteristics. We here describe the metrics (summarised in Table 1) and methods used to implement them. [^3] Two of the metrics rely on Essentia implementation. Essentia is an open-source library and tools for audio and music analysis, description and synthesis, developed in the Music Technology Group at Universitat Pompeu Fabra: https://essentia.upf.edu.
Cover Song Identification (CoverID) [19, 20, 21]: Cover song identification is a task aiming to detect whether two music recordings are based on the same composition, accounting for variations in tempo, structure, and instrumentation while keeping a similar melodic or harmonic line. CoverID relies on pitch-content features and local alignment. To obtain CoverID distance, we use the implementation available in Essentia. [^4] https://essentia.upf.edu/reference/std_CoverSongSimilarity.html A low CoverID value suggests substantial composition similarity between the two analysed music samples.
@@
Table 1: Summary of the considered music similarity metrics, indicating whether values correspond to higher or lower similarity ($\downarrow$/$\uparrow$).
@@ 3.2 Experimental Approach
For this experiment, we use an in-house dataset of 30-second audio previews from the Spotify API [^10] https://developer.spotify.com/documentation/web-api, composed of over 18,000 samples and 24 music genre classes. We focus on six music genre classes defined by Spotify API internal class labels: heavy metal, afrobeats, techno, dub, cumbia and bolero. These genres were chosen for their diverse musical compositions and elements, allowing us to examine the metrics across multiple scenarios. This selection was supported using ChatGPT, which affirmed that these genres have distinct musical characteristics.
@@ 4.1 Analysing metric behaviour
Figures 2, 3, 4, 5 and 6 depict the average $\mu$ and standard deviations $\sigma$ of the different metrics per degree of replication and music genre. We observe a steady and similar behaviour by three metrics (CoverID, CLAP and DEfNet) through all studied music genres, showing higher similarity values for cases with higher replication levels (i.e., $50\%$). Standard deviation decreases with increasing replication level, which suggests less disparity within the analysed pairs. These three metrics show the sensitivity [^11] Sensitivity is understood as the capability to differentiate between degrees of replication. to estimate data replication.
Instead, KL divergence presents a different behaviour with very similar values of $\mu$ and $\sigma$ for all degrees of replication. Some sensitivity is observed in all music genres, except for dub, where the baseline mean $\mu_{b}$ is smaller than in replication cases $\mu_{r}$, despite the standard deviation being higher ($\mu_{b}$=0.757, $\sigma_{b}$=0.511; $\mu_{r}$=0.862, $\sigma_{r}$=0.462). Thus, KL divergence demonstrates the capability of detecting replication but is ineffective in distinguishing between degrees of replication.
@@ 6 Discussion and Conclusions
This work focused on validating the use of music similarity metrics for assessing data replication in AI-generated music. We hypothesise that similarity metrics are effective in estimating data replication. Therefore, we framed the scope of our study to exact data replication in music samples, while conducting a controlled forced-replication experiment with synthetic data.
@@ 6.1 Limitations and Future Work
7 Acknowledgments
The authors would like to thank Gaëtan Hadjeres and William Thong from Sony AI, as well as Dmitry Bogdanov and Pablo Alonso-Jiménez from the Music Technology Group at Universitat Pompeu Fabra, for their insightful discussions and valuable feedback throughout the development of this research.